OpenStudy (anonymous):

What is the solution to the equation? (x+8)-14=-15 so x =______

OpenStudy (science_ally):

so x is going to be a negative number right off the bat

OpenStudy (anonymous):

it is -9 right?

OpenStudy (anonymous):

Frank is in an airplane flying at 14,080 ft above sea level. He wants to land the airplane on a runway that is 22 ft below sea level. What was the total vertical distance between his current height and the runway? ft_____ can you help me with this one?

OpenStudy (science_ally):

hold on a sec or so plz

OpenStudy (anonymous):

ok

OpenStudy (science_ally):

yes -9 is correct

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

The variable t represents the temperature at 4 p.m. Which equation would best model the word problem below? The temperature at the summit of Mt. Mansfield dropped 17°F between 4 p.m. and 11 p.m. The temperature at 11 p.m. was –11°F. What was the temperature at 4 p.m.? A. 17 + 11 = t B. 11 – 17 = t C. t – 17 = –11 D. 17 – t = –11

OpenStudy (science_ally):

so the sea level problem I don't think I can help with .

OpenStudy (anonymous):

np lol but thx for trying

OpenStudy (science_ally):

on the next problem your answer is C. Because you have to do 17 from t which will lead you to -11

OpenStudy (triciaal):